# GraPE[Uncaptioned image]: A Generate-Plan-Edit Framework for Compositional T2I Synthesis

### v1

@@ Abstract
In this work, we present an alternate paradigm for T2I synthesis, decomposing the task of complex multi-step generation into three steps, (a) Generate: we first generate an image using existing diffusion models (b) Plan: we make use of Multi-Modal LLMs (MLLMs) to identify the mistakes in the generated image expressed in terms of individual objects and their properties, and produce a sequence of corrective steps required in the form of an edit-plan. (c) Edit: we make use of an existing text-guided image editing models to sequentially execute our edit-plan over the generated image to get the desired image which is faithful to the original instruction. Our approach derives its strength from the fact that it is modular in nature, is training free, and can be applied over any combination of image generation and editing models. As an added contribution, we also develop a model capable of compositional editing, which further helps improve the overall accuracy of our proposed approach. Our method flexibly trades inference time compute with performance on compositional text prompts. We perform extensive experimental evaluation across 3 benchmarks and 10 T2I models including DALLE-3 and the latest – SD-3.5-Large. Our approach not only improves the performance of the SOTA models, by upto 3 points, it also reduces the performance gap between weaker and stronger models. The code is available at https://dair-iitd.github.io/GraPE/
@@ 1 Introduction
We perform a series of experiments to evaluate the efficacy of our approach over three benchmark datasets. We compare with 10 different T2I models, and show improvement in each case, with improvements up-to a significant 20+ points in some cases. We also note that we see a more significant improvement in weaker (typically smaller) generation models, resulting in narrowing of performance gap between smaller and larger SoTA models. We also perform ablations showing the value of our object-centric approach for generating edit plans. Along with this, significant boost in performance is shown when we use the compositional image editor that we develop, as part of our pipeline. We also present detailed error analysis and point to limitations of our model in our experimental analysis.
@@ 3.3 Text Guided Image Editing
T2I models based on CLIP [38], and by extension, editing models built on it are limited by CLIP’s compositional reasoning capabilities including understanding word order in text and object relations in images [59, 28, 48]. We hypothesize and show for the first time that enhancing word order understanding in the text space and training the model on high-quality object-centric and reasoning-oriented data is key to induce compositionality in image editors. For this, we introduce PixEdit, a text-guided image editing model based on PixArt-Sigma [11] and trained on the reasoning-centric dataset used by AURORA [29]. With the T5-XXL [39] language model as its text encoder, PixEdit demonstrates enhanced performance on compositional and spatially complex edits, benefiting from both robust language comprehension and reasoning-centric training data. Providing gains over all existing image-editing models for the generation task in our GraPE framework. Note that PixEdit is developed to be a general-purpose image editor, with enhanced compositional and object centric editing capabilities, and is not specific to the generation task we tackle. For additional details about PixEdit refer Supplementary Section 9.
@@ 3.4 Implementation Details
{NiceTabular}
[…]
A porcelain pot with tulips and a metal can with orchids and a glass jar with sunflowers [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Remove sunflowers from image. Add a glass jar on table. Put sunflowers in the glass jar.
Three chairs are arranged in a row. A gray duck is positioned on the right side of the rightmost chair. A skyline in background. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Remove the duck from the leftmost chair. Remove the duck from the middle chair. Position a gray duck on the right side of the rightmost chair.
Change the blue apple to a blue bowl. Change the red apple to a pink apple.
Table 1: Iterative results by applying GraPE. on images generated by SD3.5 and SDXL, these images are edited via the proposed PixEdit editing model. Please see 8 for additional details on the plans generated for these images.
Figure 3: Experimental results showcasing the maximum gain in DSG score by GraPE with both AURORA and PixEdit as editing models. The figure presents both DSG and DSG (w/o dependency) scores. The percentage gain is measured over DSG scores.
@@
Table 2: Results on Concept-mix benchmark  GraPEPixEdit
@@ 4.1 Experimental Settings
To evaluate the effectiveness of GraPE, we conduct a comprehensive assessment across 10 state-of-the-art text-to-image (T2I) models of varying sizes and capabilities. These models include: Stable Diffusion v1.5 [42], Structure Diffusion [16], Stable Diffusion V2.1 [42], LMD [31], SD-XL [37], PixArt-$\alpha$ [10], DeepFloyd IF [2], PlaygroundV2.5 [30], Dalle 3 [5], and the latest Stable Diffusion v3.5 Large [46]. This selection encompasses a diverse range of recent T2I models, allowing for a thorough evaluation of GraPE’s robustness across different architectures and configurations.
@@ 4.2 Results
GraPE[Uncaptioned image] Improves Compositional T2I Synthesis GraPE can be used with any given back-bone in a plug-n-play manner, and substantially improves performance across the board for all of the 10 T2I models tested in this study, including SoTA models like SD3.5 [46]. Results on T2I Compbench and Flickr-Bench are present in Figure 3. Results on ConceptMix are present in Table 2. On T2I Compbench we see improvements ranging from 17.6% in the case of SD1.5 to nearly 2% for the latest SD3.5 Large model. On ConceptMix as well, we see large performance gains across prompt complexities $K$=$\{1,3,5,7\}$. For prompts with higher complexity ($K$=$7$), we see gains ranging from 3.1% to 35.5% across models.
@@
GraPE[Uncaptioned image] scales well with more compute Our approach can be viewed from a lens of using extra inference time compute (calls to the editing model) for creating images better aligned with complex prompts, achieving better performance on the T2I Synthesis task. GraPE flexibly trades off performance with the amount of compute (or the number of edits at inference time). See Fig. 4 for results on four representative models. This also shows that each planning step of GraPE is important on average for improving on the T2I synthesis task gradually.
For additional results, including comparisons with other recent approaches, see Supplementary Section 8.
@@ 4.3 Ablations
This section highlights the importance of object and attribute centric decomposition of plans to generate effective editing instructions. We modify the few-shot structure of the MLLM planner to skip the decomposition into image and text elements and prompt it to identify errors and suggest editing plans given the text-prompt and generated image. We refer to the pipeline thus constructed as GraPEnaive. Table 3 presents the results of both GraPE and GraPEnaive on $K$=$7$ subset of ConceptMix benchmark for a subset of models. See Supplementary Section 7 for more ablations.
@@
Figure 4: Trend of GPT-QA score with increasing number of editing steps.
Figure 5: Average Steps per Plan for ConceptMix Benchmark across the models.
@@ 4.4 Analysis and Insights
We conduct a detailed examination of each component in GraPE, focusing specifically on identifying and analyzing failure cases in both the MLLM planner and the editing model. For this purpose, we curate a dataset of 120 image-prompt pairs, uniformly sampled across all 10 models and 6 benchmarks. This dataset is carefully evaluated by six human annotators, compensated well above the national average hourly wage.
@@
In Fig. 5, we see a gradual increase in the average number of plans as we increase the prompt complexity from $K$=$1$ to $K$=$7$ using the ConceptMix benchmark. This shows our planner is able to recognize the complexity of the prompt and generate a longer plan in response. We also see that on average, more capable T2I models like SD3.5 require fewer editing steps compared to weaker models such as SD1.5.
Additional qualitative examples of our approach are shown in Supplementary Section 10.
@@ 8 Additional and Detailed Results:
We compare the performance of GraPE against RPG [57], a recent LLM-based image generation method. RPG leverages an LLM to decompose complex image generation tasks into simpler subtasks within localized subregions and employs complementary-regional-diffusion to coherently merge the generated subregions into a unified image. The comparison is conducted on the ConceptMix benchmark, with RPG using SD-XL [37] as its base model, while GraPE employs the PixEdit editing model. Results are summarized in Table 5.
Our evaluation reveals that RPG fails to generate valid plans for over 50% of the samples in the comparison, primarily due to frequent parsing errors. For the remaining samples, GraPE demonstrates significant improvements in GPT-QA accuracy over images generated by RPG. Specifically, GraPE achieves gains of $21.5\%$, $3.0\%$, $18.0\%$, and $20.3\%$ for $K$=$1$, $K$=$3$, $K$=$5$, and $K$=$7$, respectively.
@@
Results on ConcetpMix Benchmark using GraPE with AURORA
@@
Method Concept K=1 Concept K=3 Concept K=5 Concept K=7
RPG [57] $0.696$±0.015 0.845 ±0.008 $0.694$0.005 0.715 ±0.007 $0.583$0.002 0.688 ±0.005 $0.388$0.007 0.467 ±0.005
Table 5: Comparison of RPG [57] and GraPE on ConceptMix
@@ 9.3.1 Analysis of Improvement in PixEdit
Upgraded Backbone Diffusion Model: Transitioning from SD1.5 to PixArt-Sigma.
Refined Pre-Training Data: Replacing the Instruct-Pix2Pix dataset with the real-only subset of the Seed-Edit dataset for pre-training.
@@ 10 Additional Qualitative Examples:
Table 9: Results on Concept-mix benchmark - GraPEAURORA
@@
Table 10: GPT-QA scores on T2I Comp-Bench, GraPEX refers to using GraPE with X editing model
@@
Table 11: GPT-QA scores on Flick-Bench, GraPEX refers to using GraPE with X editing model

### v2

@@ Abstract
In this work, we present an alternate paradigm for T2I synthesis, decomposing the task of complex multi-step generation into three steps, (a) Generate: we first generate an image using existing diffusion models (b) Plan: we make use of Multi-Modal LLMs (MLLMs) to identify the mistakes in the generated image expressed in terms of individual objects and their properties, and produce a sequence of corrective steps required in the form of an edit-plan. (c) Edit: we make use of an existing text-guided image editing models to sequentially execute our edit-plan over the generated image to get the desired image which is faithful to the original instruction. Our approach derives its strength from the fact that it is modular in nature, is training free, and can be applied over any combination of image generation and editing models. As an added contribution, we also develop a model capable of compositional editing, which further helps improve the overall accuracy of our proposed approach. Our method flexibly trades inference time compute with performance on compositional text prompts. We perform extensive experimental evaluation across 3 benchmarks and 13 T2I models including DALLE-3 and the latest – SD-3.5-Large. Our approach not only improves the performance of the SOTA models, by upto 3 points, it also reduces the performance gap between weaker and stronger models. The code is available at https://dair-iitd.github.io/GraPE/
@@ 1 Introduction
We perform a series of experiments to evaluate the efficacy of our approach over three benchmark datasets. We compare with 13 different T2I models, and show improvement in each case, with improvements up-to a significant 20+ points in some cases. We also note that we see a more significant improvement in weaker (typically smaller) generation models, resulting in narrowing of performance gap between smaller and larger SoTA models. We also perform ablations showing the value of our object-centric approach for generating edit plans. Along with this, significant boost in performance is shown when we use the compositional image editor that we develop, as part of our pipeline. We also present detailed error analysis and point to limitations of our model in our experimental analysis.
@@ 3.2 Multi-Modal Planner
{NiceTabular}
[…]
In the image, a corgi has a tiny apple positioned on top of its head. A cactus and a fork are placed nearby. Additionally, there is a duck with metallic texture in the scene. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Replace the cactus on the corgi’s head with a tiny apple Add a duck with metallic texture to the scene
Change the blue apple to a blue bowl Change the red apple to a pink apple
A boy wearing a black shirt and khaki pants is laying down, and is surrounded by stuffed animals. [Uncaptioned image] [Uncaptioned image]
Change the pants to khaki color
Table 1: Results produced by GraPE. on images generated by SDXL, SD3.5-L and PixArt, using PixEdit as image-editor.
@@ 3.3 Text Guided Image Editing
T2I models based on CLIP [41], and by extension, editing models built on it are limited by CLIP’s compositional reasoning capabilities including understanding word order in text and object relations in images [63, 29, 50]. We hypothesize and show for the first time that enhancing word order understanding in the text space and training the model on high-quality object-centric and reasoning-oriented data is key to induce compositionality in image editors. For this, we introduce PixEdit, a text-guided image editing model based on PixArt-Sigma [11] and trained on the reasoning-centric dataset used by AURORA [30]. With the T5-XXL [42] language model as its text encoder, PixEdit demonstrates enhanced performance on compositional and spatially complex edits, benefiting from both robust language comprehension and reasoning-centric training data. Providing gains over all existing image-editing models for the generation task in our GraPE framework. Note that PixEdit is developed to be a general-purpose image editor, with enhanced compositional and object centric editing capabilities, and is not specific to the generation task we tackle. Supp. Tables 6 and 7 provide comparison of PixEdit’s editing capability on two existing image editing benchmarks with other baselines. For detailed discussion about implementation and analysis of improvement in PixEdit refer Supp. Section 9.3.
@@ 3.4 Implementation Details
Figure 3: Experimental results showcasing the maximum gain in DSG score by GraPE with both AURORA and PixEdit as editing models. The figure presents both DSG and DSG (w/o dependency) scores. The percentage gain is measured over DSG scores. The absolute values are presented in Table 10 and 11
@@
Flux-schnell [31] $0.902$0.002 0.918 ±0.002 $0.820$0.003 0.864 ±0.001 $0.786$0.005 0.804 ±0.008 $0.775$0.004 0.779 ±0.004
RAG [12] $0.815$0.009 0.866 ±0.003 $0.668$0.007 0.718 ±0.005 $0.665$0.002 0.721 ±0.004 $0.520$0.003 0.628 ±0.001
RPG [61] $0.696$0.015 0.845 ±0.008 $0.694$0.002 0.715 ±0.007 $0.583$0.002 0.688 ±0.005 $0.388$0.007 0.467 ±0.005
Table 2: Results on Concept-mix benchmark  GraPEPixEdit. For complementary table with AURORA as editing model refer Table 9.
@@ 4.1 Experimental Settings
To evaluate the effectiveness of GraPE, we conduct a comprehensive assessment across 13 state-of-the-art text-to-image (T2I) models of varying sizes and capabilities. These models include: Stable Diffusion v1.5 [44], Structure Diffusion [17], Stable Diffusion V2.1 [44], SD-XL [40], PixArt-$\alpha$ [10], DeepFloyd IF [2], PlaygroundV2.5 [32], Dalle 3 [5], Stable Diffusion v3.5 Large [48], Flux-schnell [31] and three baselines which adopts MLLMs for complex T2I generation: LMD [33], RPG [61] and RAG [12]. We also compare GraPE with recently proposed SLD [57] in Suppl. Section 8, which is a post-hoc self correction method using MLLMs. This selection encompasses a diverse range of recent T2I models, allowing for a thorough evaluation of GraPE’s robustness across different architectures and configurations.
@@ 4.2 Results
GraPE[Uncaptioned image] Improves Compositional T2I Synthesis GraPE can be used with any given back-bone in a plug-n-play manner, and substantially improves performance across the board for all of the 13 T2I models tested in this study, including SoTA models like SD3.5 [48]. Results on T2I Compbench and Flickr-Bench are present in Figure 3. Results on ConceptMix are present in Table 2. On T2I Compbench we see improvements ranging from 17.6% in the case of SD1.5 to nearly 2% for the latest SD3.5 Large model. On ConceptMix as well, we see large performance gains across prompt complexities $K$=$\{1,3,5,7\}$. For prompts with higher complexity ($K$=$7$), we see gains ranging from 3.1% to 35.5% across models.
@@
GraPE[Uncaptioned image] scales well with more compute Our approach can be viewed from a lens of using extra inference time compute (calls to the editing model) for creating images better aligned with complex prompts, achieving better performance on the T2I Synthesis task. GraPE flexibly trades off performance with the amount of compute (or the number of edits at inference time). See Fig. 4(a) for results on four representative models. This also shows that each planning step of GraPE is important on average for improving on the T2I synthesis task gradually.
@@ 4.3 Ablations
This section highlights the importance of object and attribute centric decomposition of plans to generate effective editing instructions. We modify the few-shot structure of the MLLM planner to skip the decomposition into image and text elements and prompt it to identify errors and suggest editing plans given the text-prompt and generated image. We refer to the pipeline thus constructed as GraPEnaive. Table 3 presents the results of both GraPE and GraPEnaive on $K$=$7$ subset of ConceptMix benchmark for a subset of models. See Supp. Section 7 for further details.
@@
Figure 4: (a) Trend of GPT-QA score with increasing number of editing steps. (b) Average edit steps per plan for ConceptMix benchmark across the models
@@ 4.4 Analysis and Insights
We conduct a detailed examination of each component in GraPE, focusing specifically on identifying and analyzing failure cases in both the MLLM planner and the editing model. For this purpose, we curate a dataset of 120 image-prompt pairs, uniformly sampled across all 13 models and 6 benchmarks. This dataset is carefully evaluated by six human annotators, compensated well above the national average hourly wage.
@@
In Fig. 4(b), we see a gradual increase in the average number of plans as we increase the prompt complexity from $K$=$1$ to $K$=$7$ using the ConceptMix benchmark. This shows our planner is able to recognize the complexity of the prompt and generate a longer plan in response. We also see that on average, more capable T2I models like SD3.5 require fewer editing steps compared to weaker models such as SD1.5.
For inference cost analysis and runtime comparison refer to Suppl. Section 8. Additional qualitative examples of our approach are shown in Suppl. Section 10.
@@ 8 Additional and Detailed Results:
In this section, we extend our comparison to recently proposed SLD framework [57] that generates an image from the input prompt, assesses its alignment with the prompt, and performs self-corrections on the inaccuracies in the generated image. Unlike GraPE, SLD operates in image layout-space which is either generated using either LLMs or open-vocabulary object detectors. The corrective steps are defined over object latents within the layout. Table 5 shows the performance of SLD when compared to GraPE  on ConceptMix benchmark. GraPE  consistently outperforms SLD in rectifying generated images across models and prompt-complexities.
Note on RPG: We find that RPG fails to generate valid plans for over 50% of the samples in the comparison, primarily due to frequent parsing errors, therefore the numbers reported are on the successfully generated samples only.
Complexity $\rightarrow$ K=1 K=3 K=5 K=7
Method $\rightarrow$ Base SLD GraPE Base SLD GraPE Base SLD GraPE Base SLD GraPE
SD1.5 0.81 0.87 0.89 0.61 0.64 0.75 0.50 0.61 0.69 0.45 0.55 0.61
StructDiff 0.82 0.90 0.90 0.61 0.66 0.72 0.54 0.64 0.70 0.44 0.51 0.57
SD2.1 0.83 0.87 0.87 0.64 0.68 0.74 0.58 0.63 0.69 0.47 0.57 0.63
SDXL 0.85 0.85 0.88 0.71 0.68 0.78 0.64 0.64 0.73 0.52 0.53 0.63
PixArt 0.81 0.87 0.87 0.67 0.68 0.72 0.65 0.67 0.74 0.51 0.55 0.63
DeepFloyd-IF 0.88 0.91 0.92 0.68 0.71 0.77 0.66 0.68 0.75 0.58 0.62 0.66
PlaygroundV2.5 0.91 0.91 0.96 0.74 0.73 0.79 0.66 0.66 0.72 0.54 0.54 0.64
Dalle3 0.95 0.89 0.95 0.83 0.77 0.86 0.81 0.77 0.83 0.73 0.65 0.74
SD3.5 Large 0.93 0.90 0.95 0.81 0.80 0.82 0.80 0.73 0.83 0.76 0.72 0.78
Flux 0.90 0.90 0.92 0.82 0.82 0.87 0.79 0.73 0.81 0.78 0.68 0.78
LMD 0.86 0.85 0.87 0.71 0.71 0.77 0.64 0.64 0.73 0.59 0.59 0.67
RAG 0.81 0.82 0.87 0.67 0.71 0.72 0.66 0.68 0.72 0.52 0.56 0.63
RPG 0.70 0.73 0.85 0.69 0.68 0.72 0.58 0.64 0.69 0.39 0.54 0.47
Table 5: Comparison of performance between SLD and GraPE  when applied over Base images as post-hoc correction.
@@
Results on ConceptMix Benchmark using GraPE with AURORA
@@
Cost-Analysis
We quantify the additional cost incurred by GraPE primarily through the planning phase, which leverages GPT-4o for generating precise, atomic edit instructions. Our measurements indicate that this planning stage costs approximately \$0.013 per image, amounting to roughly \$52 for processing the entire ConceptMix benchmark. We note that knowledge distillation methods described in OmniEdit [54] and LLaVa [35] could help by transferring GPT-4o level capabilities to smaller open-source models, potentially reducing inference costs. The modular design of the framework enables a flexible trade-off between performance and compute cost.
Runtime-Comparison
We carry out runtime evaluation of GraPE on an A100 GPU. We find that the planning phase requires an average of 20 seconds per image, with the editing phase adding approximately 3 seconds for each atomic edits—given that we perform an average of 2.53 edit steps, calculated across benchmarks and models. This is in contrast to other post-hoc correction methods, such as SLD and RAG, which average at 52.5 and 55 seconds per image, respectively. Considering that the average base generation time for the top 5 performing text-to-image models is 20 seconds (approx), our approach effectively leverages iterative refinement to bridge the performance gap between weaker and stronger models without incurring excessive computational delays.
@@ 9.3.1 Analysis of Improvement in PixEdit
(1) Upgraded Backbone Diffusion Model: Transitioning from SD1.5 to PixArt-Sigma.
(2) Refined Pre-Training Data: Replacing the Instruct-Pix2Pix dataset with the real-only subset of the Seed-Edit dataset for pre-training.
@@ 10 Additional Qualitative Examples:
Flux-schnell [31] $0.902$0.002 0.940 ±0.004 $0.820$0.003 0.856 ±0.001 $0.786$0.005 0.781 ±0.002 $0.775$0.004 0.748 ±0.007
RAG [12] $0.815$0.009 0.865 ±0.005 $0.668$0.007 0.718 ±0.005 $0.665$0.002 0.687 ±0.007 $0.520$0.003 0.604 ±0.004
RPG [61] $0.696$0.015 0.831 ±0.001 $0.694$0.002 0.756 ±0.005 $0.583$0.002 0.688 ±0.007 $0.388$0.007 0.505 ±0.003
Table 9: Results on Concept-mix benchmark - GraPEAURORA, Complementary to Table 2
@@
Flux-Schnell [31] $0.901$0.003 $0.892$0.004 $0.915$0.006 $0.901$0.007 0.914 ±0.003 0.902 ±0.003
RAG [31] $0.822$0.003 $0.813$0.004 $0.867$0.006 $0.852$0.007 0.853 ±0.003 0.834 ±0.003
Table 10: GPT-QA scores on T2I Comp-Bench presented in Fig 3, GraPEX refers to using GraPE with X editing model
@@
Flux-Schnell [31] $0.884$0.003 $0.864$0.004 $0.876$0.003 $0.846$0.004 0.890 ±0.001 0.871 ±0.002
RAG [31] $0.750$0.003 $0.676$0.004 $0.809$0.003 $0.756$0.004 0.816 ±0.002 0.782 ±0.002
Table 11: GPT-QA scores on Flick-Bench presented in Fig 3, GraPEX refers to using GraPE with X editing model